Calculate the missing information for the installment loan that is being paid off early.

Number of
Payments
Payments
Made
Payments
Remaining
Sum-of-the-
Digits Payments
Remaining
Sum-of-the-
Digits Number
of Payments
Rebate
Fraction
36 21
Answer #1

Solution :-

(i)Payments Remaining = Total payments - Payments Made

= 36 - 21 = 15 Payments Remaining

Sum of digits is to be calculated by n(n+1) /2

(ii) Sum of digits Payments Remaining (SOD Payment Remaining) = 15*(15+1) / 2 = 120

(iii) Sum of digits Number of Payments (SOD Number of Payments) = 36*(36+1) / 2 = 666

(iv) Rebate Fraction = SOD Payment remaining / SOD Number of Payment = 120 / 666 = 18.02%